4x 6 -24x 5 +20x 4 +68x 3 -44x 2 -40x+k is a perfect square find k.

4x6-24x5+20x4+68x3-44x2-40x+k is a perfect square


find k.


 

Grade:10

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
As given it is a perfect square
Let given expression =(ax3-bx2+cx+d)2
Expanding above expression we get
=(a2x6-2abx5+(b2+2ac)x4+(2ad-2bc)x3+(c2-2bd)x2+2cdx+d2)
By comparing above expression with given expression we get
a2=4 so a=2
-2ab=-24 so b=6
b2+2ac=20 so c=-4
2ad-2bc=68 so d=5
Asked to find k which is d2=25
So value of k=25